Please help me figure out the problem below.
Mark invests $8000 in an account that pays 12% interest and $2000 in one that pays 8%. If he leaves the money in the accounts for the same length of time, how long must he leave it to gain $5600 in interest?
So far I have: P= 8000 R=.12 T=X I= 960X
* *
P= 2000 R= .08 T=Y I= 160Y
HELP ME PLEASE!

Answer by mananth(16949)

Interest on I + Interest on II = 5600
I continue where you left
960x+160y=5600
but x=y
960x+160x=5600
1120x=5600
/1120
x= 5
5 years